If my age now plus my age 6 years ago is 24,how old will I be in 2 years?
Answer by alicealc(293)

x + (x - 6) = 24
x + x - 6 = 24
2x - 6 = 24
2x = 24 + 6
2x = 30
x = 30/2 = 15
2 years later = x + 2 = 15 + 2 = 17

so, two years later I will be 17 years old